# Mean-field Underdamped Langevin Dynamics and its Space-Time Discretization

### 26 Dec 2023 v1

@@ 1 Introduction
    We sharpen the convergence guarantees for both the mean-field underdamped Langevin dynamics and its space-discretization under the same set of assumptions utilized by Chen et al. (2023) (summarized in Theorems 1 and 2). In particular, we sharpen the dependence on the smoothness constant.
@@ 2.3 Definitions and assumptions
Assumption 2 ($\mathscr{L}$-smoothness).
@@
for any $\mu_{1},\mu_{2}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $x_{1},x_{2}\in\mathbb{R}^{d}$. We assume $\mathscr{L}\geq 1$ throughout this paper.
Assumption 3 (LSI).
@@
Note that if $\hat{\mu}^{x}(x)\propto\exp(-\frac{\delta F}{\delta\mu}(\mu^{x},x))$ satisfies LSI with constant $\rho>0$, Assumption 3 is satisfied with the choice $\mathscr{C}_{\textsf{LSI}}=\frac{1}{2}\land\rho$. And we refer our readers to Chen et al. (2023) and Suzuki et al. (2023) for the verification of Assumptions 1-3 in a variety of settings, including the neural network setting and density estimation setting. Beyond Assumptions 1-3, we introduce four additional assumptions that will be required to show the global convergence of our space-time discretization
Assumption 4 (Bounded Gradient).
@@
Notably, Suzuki et al. (2023) assume that $F$ can be decomposed as $F(\mu)=U(\mu)+\mathbb{E}_{x\sim\mu}[r(x)]$ where $\|D_{\mu}U(\mu,x)\|\leq R$ for any $\mu\in\mathcal{P}(\mathbb{R}^{d})$, $x\in\mathbb{R}^{d}$, and where $r(x)$ is $\lambda_{2}$-smooth with $\nabla r(0)=0$ in order to establish the convergence of their space-time discretization of MLD. Thus, their assumption, $\|D_{\mu}F(\mu,x)\|\leq\|D_{\mu}U(\mu,x)\|+\|\nabla r(x)\|\leq R+\lambda_{2}\|x\|$, implies Assumption 4 with the choice $\mathscr{L}\geq\max\{R,\lambda_{2}\}$.
Assumption 5.
For all $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the proximal Gibbs distribution (5) satisfies $\mathbb{E}_{\hat{\mu}}\|\cdot\|^{2}=O(d)$
Assumption 6.
@@
Assumptions 5-7 are needed for bounding the second moment of the iterates $(x_{t},v_{t})$ and $(x_{t}^{i},v_{t}^{i})$ along MULD and NULD, which is crucial for the establishment of our discrete-time convergence. While these assumptions are sufficient, they may not be necessary for the iterates to be bounded. Nevertheless, we argue these assumptions are not too restrictive by providing several examples satisfying them in Section 4.
@@ 2.4 Related work
Techniques for establishing the continuous-time convergence of the mean-field underdamped systems and their space-discretization (N-particle systems) are centered around coupling and hypocoercivity, also known as functional approaches (Villani, 2009). The coupling approach generally constructs a joint probability of the mean-field and N-particle systems to make the analytic comparison between them. Based on coupling approaches, Guillin et al. (2022); Bolley et al. (2010); Bou-Rabee and Schuh (2023) obtain the continuous-time convergence of the underdamped systems with mean-field interaction and its space-discretization. Duong and Tugaut (2018); Kazeykina et al. (2020) study the ergodicity of the MULD without a quantitative rate. Under the setting of small mean-field dependence, Kazeykina et al. (2020) obtain the exponential contraction using coupling techniques in Eberle et al. (2019a, b). However, the small mean-field dependence is not satisfied in many applications. The functional approach (hypocoercivity) generally constructs appropriate Lyapunov functionals and studies how their values change along the dynamics. Based on hypocoercivity, Monmarché (2017); Guillin et al. (2021); Guillin and Monmarché (2021); Bayraktar et al. (2022) establish the exponential convergence of the mean-field underdamped systems and its propagation of chaos by constructing a suitable Lyapunov functional. Nevertheless, most of the works above only consider specific settings of MULD such as singular interactions and two-body interactions, which restricts the application to real-world problems. Setting $\gamma=1$, Chen et al. (2023) establish the exponential convergence of MULD and NULD using the hypocoercivity technique in Villani (2009). Under Assumptions 1-3, they derive the convergence without restricting the size of interactions, which subsumes many settings above. Notably, the techniques of our Theorems 1 and 2 are adopted from Chen et al. (2023) based on hypocoercivity where we consider other choices of $\gamma$ to improve the decaying rate of MULD and NULD established in Chen et al. (2023).
@@ 3.1 Convergence analysis
Both proofs utilize a Lyapunov functional similar to the one introduced by Chen et al. (2023). In particular, Theorem 1 is established by showing the following functional,
@@
is decaying along the trajectory of the dynamics. Our second Theorem 2 establishes convergence of the N-particle system NULD. Using the notation $\textbf{x}=(x^{1},...,x^{N})$, $\textbf{v}=(v^{1},...,v^{N})$ and $\mu^{N}=\otimes_{i=1}^{N}\mu^{i}$, we take $\mu_{*}^{N}$ to be the limiting distribution of NULD satisfying $\mu_{*}^{N}\propto\exp\left(-NF(\mu_{\textbf{x}})-\frac{1}{2}\|\textbf{v}\|^{2}\right)$. With these definitions, we obtain our guarantee by showing the functional,
@@
$\mathcal{F}^{N}(\mu)$ $\triangleq\int NF(\mu_{\textbf{x}})+\frac{1}{2}\|v\|^{2}\mu^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})+\text{Ent}(\mu^{N}),$
@@
Theorem 1 (Mean-field underdamped Langevin dynamics).
If Assumptions 1-3 hold, $\mu_{0}$ has finite second moment, finite entropy and finite Fisher information, then the law $\mu_{t}$ of the MULD with $\gamma=\sqrt{\mathscr{L}}$ and $\mathcal{E}$ defined in (8) satisfy,
@@
Theorem 2 (N-particle underdamped Langevin dynamics).
If Assumptions 1-3 hold, $\mu_{0}^{N}$ has finite second moment, finite entropy, finite Fisher information, and $N\geq(\mathscr{L}/\mathscr{C}_{\textsf{\em LSI}})\left(32+24\mathscr{L}/\mathscr{C}_{\textsf{\em LSI}}\right)$, then the joint law $\mu_{t}^{N}$ of the NULD with $\gamma=\sqrt{\mathscr{L}}$ and $\mathcal{E}^{N}$ defined in (9) satisfy
@@
Theorem 2 implies the non-uniform-in-N convergence of NULD, which incorporates a bias term involving $N$ due to the particle approximation. Our proof technique mirrors that of Chen et al. (2023), however, we demonstrate that a faster convergence and smaller bias in Theorem 2 can be achieved by choosing $\gamma=\sqrt{\mathscr{L}}$ instead of $\gamma=1$ (see Table 2).
@@
Theorem 3 (Mean-field underdamped Langevin algorithm).
In addition to the assumptions specified in Theorems 1, let Assumptions 4-6 hold. Denote $\bar{\mu}_{K}$ the law of $(x_{K},v_{K})$ of the MULA and $\kappa\triangleq\mathscr{L}/\mathscr{C}_{\sf{LSI}}$. Then in order to ensure $\|\bar{\mu}_{K}-\mu_{*}\|_{\textsf{\em TV}}\leq\epsilon$, it suffices to choose $\gamma=\sqrt{\mathscr{L}}$, $\bar{\mu}_{0}=\mathcal{N}(0,I_{2d})$, and
@@
Theorem 4 (N-particle underdamped Langevin algorithm).
In addition to the assumptions specified in Theorem 2, let Assumptions 4, 5 and 7 hold. Denote $\bar{\mu}^{i}_{K}$ the law of $(x_{K}^{i},v_{K}^{i})$ of the NULA for $i=1,...,N$ and $\kappa\triangleq\mathscr{L}/\mathscr{C}_{\mathsf{LSI}}$. Then in order to ensure $\frac{1}{N}\sum_{i=1}^{N}\|\bar{\mu}_{K}^{i}-\mu_{*}\|_{\textsf{\em TV}}\leq\epsilon$, it suffices to choose $\gamma=\sqrt{\mathscr{L}}$, $\bar{\mu}^{N}_{0}=\mathcal{N}(0,I_{2Nd})$,
@@ 3.2 Proof sketches
For the continuous-time results, we outline the proof of Theorem 1 (and analogously Theorem 2) in this section to provide intuition for how choosing $\gamma=\sqrt{\mathscr{L}}$ can improve the decaying rate of MULD. We begin with a review of some notations of hypocoercivity in Villani (2009); Chen et al. (2023):
@@
where $u_{t}=\log\frac{\mu_{t}}{\hat{\mu}_{t}}.$ Inheriting the analysis of Theorem 2.1 in Chen et al. (2023) and Lemma 32 in Villani (2009), we show that for a general $\gamma$, the Lyapunov functional (8) is decreasing along MULD satisfying
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}(\mu_{t})\leq-Y_{t}^{\mathsf{T}}\mathcal{K}Y_{t}\leq-Y_{t}^{\mathsf{T}}Y_{t}$ $\leq-\mathscr{C}_{\textsf{LSI}}(\mathcal{F}(\mu_{t})-\mathcal{F}(\mu_{*}))-\frac{1}{4\mathscr{L}+2}\mathsf{FI}_{S}(\mu_{t}\|\hat{\mu_{t}})$
@@
Applying Grönwall’s inequality leads to the decaying rate $O(\mathscr{C}_{\textsf{LSI}}/\mathscr{L})$ of MULD ($\gamma=1$) in Chen et al. (2023). If we specify $\gamma=\sqrt{\mathscr{L}}$, we can choose $b=1/\sqrt{\mathscr{L}},\,a=2,\,c=1/\mathscr{L}$. Then we obtain $\lambda_{\textsf{min}}(\mathcal{K})=2/\sqrt{\mathscr{L}}$ and
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}(\mu_{t})\leq-Y_{t}^{\mathsf{T}}\mathcal{K}Y_{t}\leq-\frac{2}{\sqrt{\mathscr{L}}}Y_{t}^{\mathsf{T}}Y_{t}$ $\leq-\frac{2\mathscr{C}_{\textsf{LSI}}}{\sqrt{\mathscr{L}}}(\mathcal{F}(\mu_{t})-\mathcal{F}(\mu_{*}))-\frac{1}{3\sqrt{\mathscr{L}}}\mathsf{FI}_{S}(\mu_{t}\|\hat{\mu}_{t})$
@@
Applying Grönwall’s inequality leads to the improved decaying rate $O(\mathscr{C}_{\textsf{LSI}}/\sqrt{\mathscr{L}})$ of MULD ($\gamma=\sqrt{\mathscr{L}}$) in our Theorem 1. We defer the whole proof to Appendix C.
For discretization errors, we outline the proof of Theorem 3 (and analogously Theorem 4) in this section. Let $(\mu_{t})_{t\geq 0}$ and $(\bar{\mu}_{t/h})_{t\geq 0}$ represent the law of MULD and MULA initialized at $\mu_{0}$. Let $\textbf{Q}_{kh}$ and $\textbf{P}_{kh}$ denote probability measures of MULD and MULA on the space of paths $C([0,kh],\mathbb{R}^{2d})$. Invoking Girsanov’s theorem (Girsanov, 1960; Oksendal, 2013; Le Gall, 2016) and Assumption 2, we can upper bound the pathwise divergence between MULD and MULA in KL divergence for stepsize $h$ and $k=1,...,K$ under Assumptions 2 and 4:
@@
where $\delta_{0}$ is Dirac measure on $\textsf{0}\in\mathbb{R}^{2d}$, and II is the second moment of $\mu_{*}$ denoted by $\textbf{m}_{2}^{2}$. Now we need to upper bound I. Under Assumption 3, $\mu_{*}$ satisfies LSI implying Talagrand’s inequality: $\textsf{I}\lesssim\textsf{KL}(\mu_{t}\|\mu_{*})/\mathscr{C}_{\textsf{LSI}}$. Under Assumptions 1 and 2, Lemma 4.2 in Chen et al. (2023) establishes the following relation between KL divergence and energy gap:
@@
where the last inequality follows from the assumption that $\mathcal{F}(\mu_{*})\geq 0$. Therefore, under Assumptions 5 on $\textbf{m}_{2}^{2}$ and 6 on $F(\mu_{0})$, our Lemma 7 establishes the upper bound of $\mathbb{E}_{\textbf{Q}_{T}}\|(x_{t},v_{t})\|^{2}$ in terms of $\mathscr{L},\,\mathscr{C}_{\textsf{LSI}}$ and $d$, which implies the uniform upper bound of $\textsf{KL}(\mu_{T}\|\bar{\mu}_{K})$. Applying Pinsker’s inequality
@@
we can convert the discretization error bound in KL divergence to that in TV distance. Combining Pinsker’s inequality and relation (12), we derive the continuous-time convergence of MULD in Theorem 1 in TV distance:
@@
$\|\bar{\mu}_{K}-\mu_{*}\|_{\textsf{TV}}$ $\leq\|\bar{\mu}_{K}-\mu_{T}\|_{\textsf{TV}}+\|\mu_{T}-\mu_{*}\|_{\textsf{TV}}\lesssim\mathcal{B}+\mathcal{V},$
where $\mathcal{V}$ vanishes exponentially fast as $T\rightarrow\infty$ and $\mathcal{B}$ is a vanishing bias as $h\rightarrow 0$. To ensure $\mathcal{V}+\mathcal{B}\leq\epsilon$, it suffices to choose $T=\widetilde{\Theta}(\sqrt{\mathscr{L}}/\mathscr{C}_{\textsf{LSI}})$ and specify $h,\,K$ as in Theorem 3. The whole proof is deferred to Appendix D.

### 01 Jan 2024 v2

@@ 1 Introduction
    We sharpen the convergence guarantees for both the mean-field underdamped Langevin dynamics and its space-discretization under the same set of assumptions utilized by Chen et al. (2023) (summarized in Theorems 3.1 and 3.2). In particular, we sharpen the dependence on the smoothness constant.
@@ 2.3 Definitions and assumptions
Assumption 2.2 ($\mathscr{L}$-smoothness).
@@
for any $\mu_{1},\mu_{2}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $x_{1},x_{2}\in\mathbb{R}^{d}$. We assume $1\leq\mathscr{L}<\infty$ throughout this paper.
Assumption 2.3 (LSI).
@@
Note that if $\hat{\mu}^{x}(x)\propto\exp(-\frac{\delta F}{\delta\mu}(\mu^{x},x))$ satisfies LSI with constant $\rho>0$, Assumption 2.3 is satisfied with the choice $\mathscr{C}_{\textsf{LSI}}=\frac{1}{2}\land\rho$. And we refer our readers to Chen et al. (2023) and Suzuki et al. (2023) for the verification of Assumptions 2.1-2.3 in a variety of settings, including the neural network setting and density estimation setting. Beyond Assumptions 2.1-2.3, we introduce four additional assumptions that will be required to show the global convergence of our space-time discretization.
Assumption 2.4 (Bounded Gradient).
@@
Notably, Suzuki et al. (2023) assume that $F$ can be decomposed as $F(\mu)=U(\mu)+\mathbb{E}_{x\sim\mu}[r(x)]$ where $\|D_{\mu}U(\mu,x)\|\leq R$ for any $\mu\in\mathcal{P}(\mathbb{R}^{d})$, $x\in\mathbb{R}^{d}$, and where $r(x)$ is $\lambda_{2}$-smooth with $\nabla r(0)=0$ in order to establish the convergence of their space-time discretization of MLD. Thus, their assumption, $\|D_{\mu}F(\mu,x)\|\leq\|D_{\mu}U(\mu,x)\|+\|\nabla r(x)\|\leq R+\lambda_{2}\|x\|$, implies Assumption 2.4 with the choice $\mathscr{L}\geq\max\{R,\lambda_{2}\}$. The next three assumptions are needed for bounding the second moment of the iterates $(x_{t},v_{t})$ and $(x_{t}^{i},v_{t}^{i})$ along MULD and NULD, which is crucial for the establishment of our discrete-time convergence:
Assumption 2.5.
For all $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the proximal Gibbs distribution (5) satisfies $\mathbb{E}_{\hat{\mu}}\|\cdot\|^{2}=O(d).$
Assumption 2.6.
@@
While Assumptions 2.5-2.7 are sufficient, they may not be necessary for the iterates to be bounded. Nevertheless, we argue these assumptions are not too restrictive by providing several examples satisfying them in Section 4.
@@ 2.4 Related work
Techniques for establishing the continuous-time convergence of the mean-field underdamped systems and their space-discretization (N-particle systems) are centered around coupling and hypocoercivity, also known as functional approaches (Villani, 2009). The coupling approach generally constructs a joint probability of the mean-field and N-particle systems to make the analytic comparison between them. Based on coupling approaches, Guillin et al. (2022); Bolley et al. (2010); Bou-Rabee and Schuh (2023) obtain the continuous-time convergence of the underdamped systems with mean-field interaction and its space-discretization. Duong and Tugaut (2018); Kazeykina et al. (2020) study the ergodicity of the MULD without a quantitative rate. Under the setting of small mean-field dependence, Kazeykina et al. (2020) obtain the exponential contraction using coupling techniques in Eberle et al. (2019a, b). However, the small mean-field dependence is not satisfied in many applications. The functional approach (hypocoercivity) generally constructs appropriate Lyapunov functionals and studies how their values change along the dynamics. Based on hypocoercivity, Monmarché (2017); Guillin et al. (2021); Guillin and Monmarché (2021); Bayraktar et al. (2022) establish the exponential convergence of the mean-field underdamped systems and its propagation of chaos by constructing a suitable Lyapunov functional. Nevertheless, most of the works above only consider specific settings of MULD such as singular interactions and two-body interactions, which restricts the application to real-world problems. Setting $\gamma=1$, Chen et al. (2023) establish the exponential convergence of MULD and NULD using the hypocoercivity technique in Villani (2009). Under Assumptions 2.1-2.3, they derive the convergence without restricting the size of interactions, which subsumes many settings above. Notably, the techniques of our Theorems 3.1 and 3.2 are adopted from Chen et al. (2023) based on hypocoercivity where we consider other choices of $\gamma$ to improve the decaying rate of MULD and NULD established in Chen et al. (2023).
@@ 3.1 Convergence analysis
Both proofs utilize a Lyapunov functional similar to the one introduced by Chen et al. (2023). In particular, Theorem 3.1 is established by showing the following functional,
@@
is decaying along the trajectory of the dynamics. Our second Theorem 3.2 establishes convergence of the N-particle system NULD. Using the notation $\textbf{x}=(x^{1},...,x^{N})$, $\textbf{v}=(v^{1},...,v^{N})$ and $\mu^{N}=\otimes_{i=1}^{N}\mu^{i}$, we take $\mu_{*}^{N}$ to be the limiting distribution of NULD satisfying $\mu_{*}^{N}\propto\exp\left(-NF(\mu_{\textbf{x}})-\frac{1}{2}\|\textbf{v}\|^{2}\right)$. With these definitions, we obtain our guarantee by showing the functional,
@@
$\mathcal{F}^{N}(\mu)$ $\triangleq\int NF(\mu_{\textbf{x}})+\frac{1}{2}\|\textbf{v}\|^{2}\mu^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})+\text{Ent}(\mu^{N}),$
@@
Theorem 3.1 (Mean-field underdamped Langevin dynamics).
If Assumptions 2.1-2.3 hold, $\mu_{0}$ has finite second moment, finite entropy and finite Fisher information, then the law $\mu_{t}$ of the MULD with $\gamma=\sqrt{\mathscr{L}}$ and $\mathcal{E}$ defined in (8) satisfy,
@@
Theorem 3.2 (N-particle underdamped Langevin dynamics).
If Assumptions 2.1-2.3 hold, $\mu_{0}^{N}$ has finite second moment, finite entropy, finite Fisher information, and $N\geq(\mathscr{L}/\mathscr{C}_{\textsf{\em LSI}})\left(32+24\mathscr{L}/\mathscr{C}_{\textsf{\em LSI}}\right)$, then the joint law $\mu_{t}^{N}$ of the NULD with $\gamma=\sqrt{\mathscr{L}}$ and $\mathcal{E}^{N}$ defined in (9) satisfy
@@
Theorem 3.2 implies the non-uniform-in-N convergence of NULD, which incorporates a bias term involving $N$ due to the particle approximation. Our proof technique mirrors that of Chen et al. (2023); however, we demonstrate that a faster convergence and smaller bias in Theorem 3.2 can be achieved by choosing $\gamma=\sqrt{\mathscr{L}}$ instead of $\gamma=1$ (see Table 1).
@@
Theorem 3.3 (Mean-field underdamped Langevin algorithm).
In addition to the assumptions specified in Theorems 3.1, let Assumptions 2.4-2.6 hold. Denote $\bar{\mu}_{K}$ the law of $(x_{K},v_{K})$ of the MULA and $\kappa\triangleq\mathscr{L}/\mathscr{C}_{\sf{LSI}}$. Then in order to ensure $\|\bar{\mu}_{K}-\mu_{*}\|_{\textsf{\em TV}}\leq\epsilon$, it suffices to choose $\gamma=\sqrt{\mathscr{L}}$, $\bar{\mu}_{0}=\mathcal{N}(0,I_{2d})$, and
@@
Theorem 3.4 (N-particle underdamped Langevin algorithm).
In addition to the assumptions specified in Theorem 3.2, let Assumptions 2.4, 2.5 and 2.7 hold. Denote $\bar{\mu}^{i}_{K}$ the law of $(x_{K}^{i},v_{K}^{i})$ of the NULA for $i=1,...,N$ and $\kappa\triangleq\mathscr{L}/\mathscr{C}_{\mathsf{LSI}}$. Then in order to ensure $\frac{1}{N}\sum_{i=1}^{N}\|\bar{\mu}_{K}^{i}-\mu_{*}\|_{\textsf{\em TV}}\leq\epsilon$, it suffices to choose $\gamma=\sqrt{\mathscr{L}}$, $\bar{\mu}^{N}_{0}=\mathcal{N}(0,I_{2Nd})$,
@@ 3.2 Proof sketches
For the continuous-time results, we outline the proof of Theorem 3.1 (and analogously Theorem 3.2) in this section to provide intuition for how choosing $\gamma=\sqrt{\mathscr{L}}$ can improve the decaying rate of MULD. We begin with a review of some notations of hypocoercivity in Villani (2009); Chen et al. (2023):
@@
where $u_{t}=\log\frac{\mu_{t}}{\hat{\mu}_{t}}.$ Inheriting the analysis of Theorem 2.1 in Chen et al. (2023) and Lemma 32 in Villani (2009), we show that for a general $\gamma$, the Lyapunov functional (8) with $S=[s_{ij}]\otimes I_{d}\in\mathbb{R}^{2d\times 2d}$ is decreasing along MULD satisfying
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}(\mu_{t})\leq-\lambda_{\textsf{min}}(\mathcal{K})Y_{t}^{\mathsf{T}}Y_{t}$ $\leq-\mathscr{C}_{\textsf{LSI}}(\mathcal{F}(\mu_{t})-\mathcal{F}(\mu_{*}))-\frac{1}{2\lambda_{\textsf{max}}(S)}\mathsf{FI}_{S}(\mu_{t}\|\hat{\mu_{t}})$
@@
Applying Grönwall’s inequality leads to the decaying rate $O(\mathscr{C}_{\textsf{LSI}}/\mathscr{L})$ of MULD ($\gamma=1$) in Chen et al. (2023). If we specify $\gamma=\sqrt{\mathscr{L}}$, we can choose $b=1/\sqrt{\mathscr{L}},\,a=2,\,c=1/\mathscr{L}$ satisfying the criteria. Then we obtain $\lambda_{\textsf{min}}(\mathcal{K})=2/\sqrt{\mathscr{L}}$ and
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}(\mu_{t})\leq-\lambda_{\textsf{min}}(\mathcal{K})Y_{t}^{\mathsf{T}}Y_{t}$ $\leq-\frac{2\mathscr{C}_{\textsf{LSI}}}{\sqrt{\mathscr{L}}}(\mathcal{F}(\mu_{t})-\mathcal{F}(\mu_{*}))-\frac{1}{\lambda_{\textsf{max}}(S)\sqrt{\mathscr{L}}}\mathsf{FI}_{S}(\mu_{t}\|\hat{\mu}_{t})$
@@
Applying Grönwall’s inequality leads to the improved decaying rate $O(\mathscr{C}_{\textsf{LSI}}/\sqrt{\mathscr{L}})$ of MULD ($\gamma=\sqrt{\mathscr{L}}$) in our Theorem 3.1. We defer the whole proof to Appendix C.
For discretization errors, we outline the proof of Theorem 3.3 (and analogously Theorem 3.4) in this section. Let $(\mu_{t})_{t\geq 0}$ and $(\bar{\mu}_{t/h})_{t\geq 0}$ represent the law of MULD and MULA initialized at $\mu_{0}$. Let $\textbf{Q}_{kh}$ and $\textbf{P}_{kh}$ denote probability measures of MULD and MULA on the space of paths $C([0,kh],\mathbb{R}^{2d})$. Invoking Girsanov’s theorem (Girsanov, 1960; Kutoyants, 2004; Le Gall, 2016) and Assumption 2.2, we can upper bound the pathwise divergence between MULD and MULA in KL divergence for stepsize $h$ and $k=1,...,K$ under Assumptions 2.2 and 2.4:
@@
where $\delta_{0}$ is Dirac measure on $\textsf{0}\in\mathbb{R}^{2d}$, and II is the second moment of $\mu_{*}$ denoted by $\textbf{m}_{2}^{2}$. Now we need to upper bound I. Under Assumption 2.3, $\mu_{*}$ satisfies LSI implying Talagrand’s inequality: $\textsf{I}\lesssim\textsf{KL}(\mu_{t}\|\mu_{*})/\mathscr{C}_{\textsf{LSI}}$. Under Assumptions 2.1 and 2.2, Lemma 4.2 in Chen et al. (2023) establishes the following relation between KL divergence and energy gap:
@@
where the last inequality follows from the assumption that $\mathcal{F}(\mu_{*})\geq 0$. Therefore, under Assumptions 2.5 on $\textbf{m}_{2}^{2}$ and 2.6 on $F(\mu_{0})$, our Lemma 7 establishes the upper bound of $\mathbb{E}_{\textbf{Q}_{T}}\|(x_{t},v_{t})\|^{2}$ in terms of $\mathscr{L},\,\mathscr{C}_{\textsf{LSI}}$ and $d$, which implies the uniform upper bound of $\textsf{KL}(\mu_{T}\|\bar{\mu}_{K})$. Applying Pinsker’s inequality
@@
we can convert the discretization error bound in KL divergence to that in TV distance. Combining Pinsker’s inequality and relation (12), we derive the continuous-time convergence of MULD in Theorem 3.1 in TV distance:
@@
$\|\bar{\mu}_{K}-\mu_{*}\|_{\textsf{TV}}$ $\leq\underbrace{\|\bar{\mu}_{K}-\mu_{T}\|_{\textsf{TV}}}_{\mathcal{B}}+\underbrace{\|\mu_{T}-\mu_{*}\|_{\textsf{TV}}}_{\mathcal{V}},$
where $\mathcal{V}$ vanishes exponentially fast as $T\rightarrow\infty$ and $\mathcal{B}$ is a vanishing bias as $h\rightarrow 0$. To ensure $\mathcal{V}+\mathcal{B}\leq\epsilon$, it suffices to choose $T=\widetilde{\Theta}(\sqrt{\mathscr{L}}/\mathscr{C}_{\textsf{LSI}})$ and specify $h,\,K$ as in Theorem 3.3. The whole proof is deferred to Appendix D.